OpenStudy (anonymous):

looking for the substitution solution for this question 2x-y=4 5x-2y=10

Parth (parthkohli):

Isolate \(\rm y\) in the first equation.

OpenStudy (anonymous):

so that would be 2x=4 which is x=2 right

Parth (parthkohli):

...?

OpenStudy (anonymous):

negative wrong this is what I thought I did add 2x+5x 2y-y 10+4

OpenStudy (anonymous):

so that would be 7x-y=14

OpenStudy (anonymous):

nubeer you love me again lol

OpenStudy (anonymous):

Isolating means: 2x - y = 4 Adding y to both sides gives: 2x = y + 4 Subtracting 4 from both sides gives: 2x - 4 = y So now that you know what y is, "substitute" it in the second equation.

OpenStudy (anonymous):

so y is 2

OpenStudy (anonymous):

then on the second equation I would do 5x-2y+2y+10 + 2y

OpenStudy (anonymous):

No. Plug in y = 2x - 4 to 5x - 2y = 10 5x - 2(2x - 4) = 10

OpenStudy (anonymous):

5x-4x-4=10

OpenStudy (anonymous):

1x-4=10

OpenStudy (anonymous):

x=14

OpenStudy (anonymous):

You made a mistake there: 5x-4x >>>-4<<< =10 5x - 2(2x - 4) = 10 5x - (2*2x) - (2*-4) = 10

OpenStudy (anonymous):

5x-4x--8=10

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Yep. Can you find y now?

OpenStudy (anonymous):

isnt it 4

OpenStudy (anonymous):

muntoo

OpenStudy (anonymous):

2x-y=4 x=2, so 2(2)-y=4 Now do the rest.

OpenStudy (anonymous):

2x-y=4 ................ eqn 1 5-2y = 10 ...............eqn 2 solve simultanously x = 4+y/2 ............. eqn 3 Now sub eqn 3 into 2 5(4+y/2) - 2y = 10 20y/2 - 2y = 10 20y - 2y = 20 18y = 20 y = 20/18 thus y = 10/9 now sub y into eqn 1 x = 4 + (10/9)/2 x = (14/10)/2 x = 28/10 thus x = 14/5 so x = 14/5 y = 10/9

OpenStudy (anonymous):

4-y=4 add 4 to both sides and y=8

OpenStudy (anonymous):

@The_G.O.A.T Your second equation is wrong. http://www.wolframalpha.com/input/?i=2x-y%3D4%2C+5x-2y%3D10

OpenStudy (anonymous):

(I mean you copied the second equation from the given wrong.)

OpenStudy (anonymous):

yeah kust left out the x but it is back in further down

OpenStudy (anonymous):

just

OpenStudy (anonymous):

@bigmommatiad Adding 4 to both sides gives 4 - y = 4 4 - y + 4 = 4 + 4 Which is 4 + 4 - y = 8 8 - y = 8

OpenStudy (anonymous):

crud

OpenStudy (anonymous):

so then you would subtract the 4 and that would be 0

OpenStudy (anonymous):

so x=2 and y=0

OpenStudy (anonymous):

@The_G.O.A.T Equation 3 is wrong. 2x-y=4 2x = 4 + y x = 2 + y/2

OpenStudy (anonymous):

@bigmommatiad Yes.

OpenStudy (anonymous):

awesome can you help me with another problem

OpenStudy (anonymous):

i would not say its wrong just left it all over 2

OpenStudy (anonymous):

a barrel containes 183 gallons and they are draining at 9 gallons a minutes write equation that models the number of gallons, g, after the t minutes the model the given situation is ??

OpenStudy (anonymous):

would it be like this 183/9=??

OpenStudy (anonymous):

They want you to write an equation. You start off with 183 and lose 9 every minute. So: gallons_left = 183 - 9 * minutes